### v2

@@ 3 Technical overview
In Section 7, we extend our results to general single-commodity instances. Here there may be multiple origins, each with a fixed outflow rate, and multiple destinations; users are however indifferent between destinations, and will choose whichever destination they can reach first.
@@ 4 Convergence to steady-state
We are given a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$. Let $\lambda\in\mathbb{R}^{V}$ be the thin flow direction for configuration $(E,E^{\infty})$; $\lambda$ is unique, due to [Koc12, Theorem 6.62]. We will refer to $\lambda$ as the steady-state direction, for reasons connected to the following definition.
@@
Let $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$ be a generalized network. There exists a constant $C_{T}$ depending only on the generalized network such that for any starting point $l^{\circ}$, the equilibrium trajectory $\ell$ starting from $l^{\circ}$ reaches steady state by time $C_{T}\cdot d(l^{\circ},I_{ss})$.
@@
Let $(x,\ell)$ be any equilibrium starting from $l^{\circ}$, and let $q$ be the corresponding queueing delays.
We will need the following lemma, which says that arcs with a positive queue have a nonzero inflow rate.
Lemma 4.4.
With $\theta$ a point of differentiability, $\ell^{\prime}_{v}(\theta)>0$ for all $v\in V$. In particular, for every resetting arc $e\in E^{*}_{\theta}$ we have $x^{\prime}_{e}(\theta)>0$.
Proof.
Let $\ell^{\prime}=\ell^{\prime}(\theta)$ and $x^{\prime}=x^{\prime}(\theta)$. Defining $V_{0}\coloneqq\set{v\in V}{\ell^{\prime}_{v}=0}$ it holds that $\sum_{e\in\delta^{-}(v)}x^{\prime}_{e}=0$ for all $v\in V_{0}$ by (TF-4) . This immediately implies $t\notin V_{0}$ as $\sum_{e\in\delta^{-}(t)}x^{\prime}_{e}=u_{0}>0$.
Assume for contradiction that $V_{0}\neq\emptyset$ and let $v_{0}$ be the first node in $V_{0}$ according to a topological ordering with respect to $E^{\prime}_{\theta}$. Clearly $v_{0}\neq s$ as $\ell^{\prime}_{s}=1$. Due to (TF-3) there has to be an incoming resetting arc $e_{0}=uv_{0}$ (with $x^{\prime}_{e_{0}}=0$). As $\ell(\theta)$ is a feasible labeling there has to be an active $v_{0}$-$t$-path. In particular there has to be an active arc $e=vw$ with $v\in V_{0}$ and $w\notin V_{0}$. But since $x^{\prime}_{e}=0$ (flow conservation at $v$) this is a contradiction to (TF-3). ∎
@@
$E^{>}$$\coloneqq\Set{vw\in E\setminus E^{\infty}}{\lambda_{w}>\lambda_{v}},$
$E^{<}$$\coloneqq\Set{vw\in E\setminus E^{\infty}}{\lambda_{w}<\lambda_{v}}\text{ and}$
$E^{=}$$\coloneqq\Set{vw\in E\setminus E^{\infty}}{\lambda_{w}=\lambda_{v}}.$
@@
[…]
@@
$d_{v}(\theta)$ $=\tfrac{\ell_{v}(\theta)}{\lambda_{v}}\;\qquad\qquad\qquad\qquad\qquad\qquad\forall v\in V$ (6)
$p_{e}(\theta)$ =$=\begin{cases}\tfrac{\ell_{w}(\theta)}{\lambda_{w}}-\tfrac{\ell_{v}(\theta)}{\lambda_{v}}-\hat{\tau}_{e}+\tfrac{s_{e}(\theta)}{\lambda_{w}}\qquad&\forall e=vw\in E^{\infty}\cup E^{>}\\ \tfrac{q_{e}(\theta)}{\lambda_{w}}&\forall e=vw\in E^{=}\cup E^{<}.\end{cases}$
@@
The more crucial part of the definition is (i); (ii) will be convenient later for technical reasons. If $\ell$ has the WSS property for all $\theta^{\prime}\geq\theta$, then $\ell$ has reached steady state in the weaker sense alluded to in the introduction: all queues will grow at linear rate from time $\theta$ onwards. This follows as by Lemma 4.4, queueing arcs are always flow carrying. Moreover, for a flow carrying arc both end points are always active and thus $q_{e}^{\prime}(\theta^{\prime})=\ell^{\prime}_{w}(\theta^{\prime})-\ell^{\prime}_{v}(\theta^{\prime})=\lambda_{w}-\lambda_{v}$. For some intuition on the distinction between this and the stronger definition of reaching steady state given in Definition 4.1, consider the example in Figure 5. This has the WSS property for all times. Initially, the shortest path to $v$ (which is not active) is through $u$, and $\ell^{\prime}_{v}(\theta)=\ell^{\prime}_{u}(\theta)=2$. After the queue on $su$ has grown sufficiently large at some point $\theta^{\prime}$ however, $sv$ becomes active, and $\ell^{\prime}_{v}(\theta^{\prime})$ decreases to 1. At this point steady state has been reached.
Figure 5: In the example the difference between the WSS property and steady state is depicted. Dotted arcs are inactive, solid arcs are active. All capacities are one and $13=\tau_{sv}>\tau_{su}+\tau_{uv}=1+$ 1. In the beginning, the label of $v$ is determined by $uv$, until $sv$ becomes active when the queue on $su$ reaches length 1. From this point, the label of $v$ is determined by $sv$ (and $uv$ becomes inactive). Since $v$ is not active the equilibrium already has the WSS property on the left, but is not yet in steady-state. Steady state is reached as soon as $sv$ becomes active.
@@
    Consider the cut $Q_{z}=\{v\in V\mid\lambda_{v}<z\}$ for any $z$. Note that all arcs that leave $Q_{z}$ are in $E^{\infty}\cup E^{>}$ and all arcs that enter $Q_{z}$ are in $E^{\infty}\cup E^{<}$. Thus, on all leaving arcs $e=vw$ as well as on the entering arcs $e=vw\in E^{\infty}$ , we have $y_{e}=\nu_{e}\lambda_{w}$, while $y_{e}=0$ on $e\in E^{<}$. Now, consider the flow $x^{\prime}$ on arcs crossing the cut. On all arcs leaving $Q_{z}$ we have either $0=x_{e}^{\prime}<y_{e}$ or $x^{\prime}_{e}\leq\ell^{\prime}_{w}\nu_{e}=\lambda_{w}\nu_{e}=y_{e}$. On the other hand, on the arcs entering $Q_{z}$ we have $x_{e}^{\prime}\geq 0=y_{e}$ for $e\in E^{<}$ and $x^{\prime}_{e}=\ell^{\prime}_{w}\nu_{e}=\lambda_{w}\nu_{e}=y_{e}$ for $e\in E^{\infty}$, by the WSS property, using that every such arc is always resetting and hence lies on an active $s$-$t$-path. Since $x^{\prime}$ and $y$ are both $s$-$t$-flows and hence have the same net flow across $Q_{z}$, all inequalities must hold with equality. In particular $x_{e}^{\prime}=0$ for every $e\in E^{<}$ crossing $Q_{z}$, and every arc $e\in E^{>}$ that crosses $Q_{z}$ is flow carrying in $x^{\prime}$. In the latter case, $v$ and $w$ both certainly lie on an active $s$-$t$-path, and so $\ell^{\prime}_{w}>\ell^{\prime}_{v}$ by the WSS property, meaning that (since we are at a point of differentiability) $e\in E^{*}_{\theta}$. Thus, we showed the properties we want for all arcs crossing some $Q_{z}$. But every arc in $E^{>}$ and $E^{<}$ crosses $Q_{z}$ for some choice of $z$, and so the claim follows.
@@
    Dual feasibility has already been observed in Lemma 4.5. Primal feasibility follows straightforwardly from (i):
@@
    • For arcs $e=vw\in E^{>}\cup E^{\infty}$ we have that $e\in E^{*}_{\theta}$ and so $x_{e}^{\prime}=\ell_{w}^{\prime}\nu_{e}=\lambda_{w}\nu_{e}=\hat{\nu}_{e}$ (using that $x^{\prime}_{e}>0$ by Lemma 4.4).
@@
Immediate, since $(d(\theta),p(\theta))$ is a feasible solution to the dual by Lemma 4.5, and $\Phi(\theta)$ is precisely the objective value of this dual solution. ∎
@@
We claim that $\Phi^{\prime}\geq-\sum_{e=vw\in\tilde{E}}\hat{\nu}_{e}\cdot\left(\frac{\ell^{\prime}_{w}}{\lambda_{w}}-\frac{\ell^{\prime}_{v}}{\lambda_{v}}\right)$. This follows from comparing terms. For $ts$ and arcs in $E^{=}$, $E^{\infty}$, $E^{>}\cap E^{\prime}_{\theta}$ and $E^{<}\setminus E^{*}_{\theta}$, the contributions are identical. For $e=vw\in E^{>}\setminus E^{\prime}_{\theta}$, we have
$-\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{v}}\right)-\nu_{e}s^{\prime}_{e}=\hat{\nu}_{e}\ell^{\prime}_{v}\left(\tfrac{1}{\lambda_{v}}-\tfrac{1}{\lambda_{w}}\right)\geq 0.$
@@
$-\nu_{e}q^{\prime}_{e}=-\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{w}}\right)\geq-\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{v}}\right).$
@@
Since $\dot{x}$ is a circulation and all flow entering $S_{z}$ is always counted, it follows that the incoming flow on arcs in $\tilde{E}$ is not smaller than the outgoing flow on arcs in $\tilde{E}$, i.e.,
$\dot{x}(\delta^{-}(S_{z})\cap\tilde{E})\geq\dot{x}(\delta^{+}(S_{z})\cap\tilde{E}).$ (9)
@@
$z\cdot\hat{\nu}(\delta^{-}(S_{z})\cap\tilde{E})$ $\geq\sum_{e=vw\in\delta^{-}(S_{z})\cap\tilde{E}}\frac{\ell^{\prime}_{w}}{\lambda_{w}}\cdot\hat{\nu}_{e}$
@@
$\geq z\cdot\hat{\nu}(\delta^{+}(S_{z})\cap\tilde{E}).$
By 4.10, $\Phi^{\prime}\geq 0$.
@@
We claim that any active arc $e=vw\in E^{\prime}_{\theta}$ satisfies $\ell^{\prime}_{w}/\lambda_{w}\geq\ell^{\prime}_{v}/\lambda_{v}$. For if not, then choosing any $z\in(\ell^{\prime}_{w}/\lambda_{w},\ell^{\prime}_{v}/\lambda_{v})$ and applying 4.11, $e\in\tilde{E}$, which we have just shown is impossible. Since $\ell^{\prime}_{s}=\lambda_{s}=1$, and every node is reachable from $s$ via an active path, $\ell^{\prime}_{v}\geq\lambda_{v}$ for all $v$.
@@
By Lemma 4.9, $\Phi$ is nondecreasing, and strictly increasing at time $\theta$ unless the WSS property is satisfied at time $\theta$. Since the WSS property implies optimality by Lemma 4.7, we deduce that $\Phi$ must be strictly increasing until the moment that $\Phi(\theta)=\mathsf{OPT}$, at which point the WSS property is satisfied for all larger times.
To bound this time (and see that it is finite), Note that $\Phi^{\prime}(\theta)$ takes on only a finite set of values depending only on the network, since its value is fully determined by the configuration $(E^{\prime}_{\theta},E^{*}_{\theta})$. More concrete, for every possible valid configuration of active and resetting edges the $\ell^{\prime}$ values are determined by the thin flow conditions and thus only depend on the capacities. So we can choose $C_{1}$ depending only on the network such that $1/C_{1}\leq\Phi^{\prime}(\theta)$ for all points of differentiability $\theta$ where $\Phi^{\prime}(\theta)>0$. So $\Phi(\theta)=\mathsf{OPT}$ for all $\theta\geq C_{1}(\mathsf{OPT}-\Phi(0))$, as desired. ∎
We remark that it follows from this that if $\Phi(\theta)=\mathsf{OPT}$, then $\ell$ satisfies the WSS property. So while we stated Lemma 4.7 (ii) as an implication, the WSS property is in fact a characterization of optimality.
@@
Consider any point of differentiability $\theta$ for which $\Phi(\theta)=\mathsf{OPT}$. Then for any $v$, $\gamma_{v}(\theta)=0$ if and only if $\ell^{\prime}_{v}(\theta)=\lambda$.
@@
First remember, as $\Phi(\theta)=\mathsf{OPT}$ the WSS property is satisfied.
Consider any active node $v\neq s$, and let $e=uv$ be any active arc entering $v$. Then $\ell^{\prime}_{v}(\theta)=\lambda_{v}$ and $\ell^{\prime}_{u}(\theta)=\lambda_{u}$ by the WSS property. Since $e$ is active and $\theta$ is a point of differentiability, either $\ell^{\prime}_{v}(\theta)=\ell^{\prime}_{u}(\theta)$, or $e\in E^{*}_{\theta}$ (meaning $x^{\prime}_{e}(\theta)>0)$ and so $e\notin E^{<}$ by Lemma 4.7. Either way, $e\in E\setminus E^{<}$. It follows that $\gamma_{v}(\theta)=0$ for all active nodes $v$.
Consider now any inactive node $v$. Let $P$ be a path of active arcs starting from an active node $u$, and with all other nodes of $P$ being inactive. If there are multiple choices for $P$, choose one with the fewest number of arcs in $E^{<}$. No arc of $P$ is resetting (since (end) nodes of resetting arcs are always active). By Lemma 4.7, this further implies that no arc of $P$ is in $E^{>}\cup E^{\infty}$, so $\lambda_{u}\leq\lambda_{v}$. Since $\theta$ is a point of differentiability, it also follows that $\ell^{\prime}_{u}(\theta)=\ell^{\prime}_{v}(\theta)$.
If $\gamma_{v}(\theta)=0$, then $P$ contains only arcs of $E^{=}$ (since there must be an active path from $s$ to $v$ in $(V,E\setminus E^{<})$, and hence such a path from the set of active nodes, and $P\subseteq(E^{<}\cup E^{=}$). So $\ell^{\prime}_{v}(\theta)=\lambda_{v}$. Conversely, if $\ell^{\prime}_{v}(\theta)=\lambda_{v}$, so that $\lambda_{u}=\ell^{\prime}_{u}(\theta)=\ell^{\prime}_{v}(\theta)=\lambda_{v}$, then $P\subseteq E^{=}$, and $\gamma_{v}(\theta)=\gamma_{u}(\theta)=0$. ∎
@@
We show that for any $v\in V\setminus S$, $\gamma^{\prime}_{v}(\theta)\leq-\Delta^{-1}<0$ (or if $\Delta=0$, then $S=V$). This suffices: after time $T_{2}=\Delta\cdot\max_{v\in V}\gamma_{v}(T_{1})$ has passed beyond time $T_{1}$, the slack $\gamma_{v}(\theta)$ has decreased to 00 for every $v\in V$, and hence all nodes are in $S$.
Consider any $v$ with $\gamma_{v}(\theta)>0$; note that $v\notin S$ by Lemma 4.13. Let $P_{1}$ be a $u$-$v$ path in $E^{=}\setminus E[S]$ with $u\in S$ and minimal slack, and moreover let $P_{1}$ be such that this property remains true for some small interval of time. That is, $\sum_{e\in P_{1}}s_{e}(\xi)=\gamma_{v}(\xi)$ for all $\xi\in[\theta,\theta+\varepsilon)$ for some small $\varepsilon>0$. Such a path must exist, given that $\theta$ is a point of differentiability; we can choose $\varepsilon$ so that all slacks are linear in this interval. Let $P_{2}$ be a $w$-$v$ path in $E^{\prime}_{\theta}\setminus E[S]$ with $w\in S$. We have that $\lambda_{v}=\lambda_{u}=\ell^{\prime}_{u}$ and $\lambda_{v}<\lambda_{w}=\ell^{\prime}_{w}=\ell^{\prime}_{v}$. The inequality follows since $P_{2}$ must use at least one arc from $E^{<}$, or otherwise $P_{2}$ would be a no slack path in $E\setminus E^{<}$. This observation further implies that if $\Delta=0$, meaning $E^{<}=\emptyset$, then $S=V$.

### v3

@@ 3 Technical overview
In Section 7, we extend our results to general single-commodity instances. Here there may be multiple origins, each with a fixed inflow rate, and multiple destinations; users are however indifferent between destinations, and will choose whichever destination they can reach first.
@@ 4 Convergence to steady-state
We are given a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$. Let $\lambda$ denote the thin flow direction for configuration $(E,E^{\infty})$; we will refer to this as the steady-state direction.
@@
Let $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$ be a generalized network. There exists a constant $C_{T}$ depending only on the generalized network such that for any starting point $l^{\circ}{\in\Omega}$, an equilibrium trajectory $\ell$ starting from $l^{\circ}$ reaches steady state by time $C_{T}\cdot d(l^{\circ},I_{ss})$.
@@
Let $(\ell,x)$ be any equilibrium starting from $l^{\circ}$, and let $q$ be the corresponding queueing delays.
@@
[…]
@@
$\text{s.t.}\qquad d_{w}-d_{v}-p_{e}$ $\leq\hat{\tau}_{e}\phantom{0}\qquad\text{for all }e=vw\in E\setminus E^{<}$
@@
$d_{v}(\theta)$ $=\tfrac{\ell_{v}(\theta)}{\lambda_{v}}\;\qquad\qquad\qquad\qquad\qquad\qquad{\text{for all }}v\in V$ (6)
[…]
@@
The more crucial part of the definition is (i); (ii) will be convenient later for technical reasons. If $\ell$ has the WSS property for all $\theta^{\prime}\geq\theta$, then $\ell$ has reached steady state in the weaker sense alluded to in the introduction: all queues will grow at linear rate from time $\theta$ onwards. This follows as by Section 2.2, queueing arcs are always flow carrying. Moreover, for a flow carrying arc both end points are always active and thus $q_{e}^{\prime}(\theta^{\prime})=\ell^{\prime}_{w}(\theta^{\prime})-\ell^{\prime}_{v}(\theta^{\prime})=\lambda_{w}-\lambda_{v}$. For some intuition on the distinction between this and the stronger definition of reaching steady state given in Section 4, consider the example in Figure 5. This has the WSS property for all times. Initially, the shortest path to $v$ (which is not active) is through $u$, and $\ell^{\prime}_{v}(\theta)=\ell^{\prime}_{u}(\theta)=2$. After the queue on $su$ has grown sufficiently large at some point $\theta^{\prime}$ however, $sv$ becomes active, and $\ell^{\prime}_{v}(\theta^{\prime})$ decreases to 1. At this point steady state has been reached.
Figure 5: In the example the difference between the WSS property and steady state is depicted. Dotted arcs are inactive, solid arcs are active. All capacities are one and $3=\tau_{sv}>\tau_{su}+\tau_{uv}=1+1$. In the beginning, the label of $v$ is determined by $uv$, until svsv becomes active when the queue on susu reaches length 1. From this point, the label of $v$ is determined by svsv (and uvuv becomes inactive). Since $v$ is not active the equilibrium already has the WSS property on the left, but is not yet in steady-state. Steady state is reached as soon as svsv becomes active.
@@
    Consider the cut $Q_{z}=\{v\in V\mid\lambda_{v}<z\}$ for any $z$. Note that all arcs that leave $Q_{z}$ are in $E^{\infty}\cup E^{>}$ and all arcs that enter $Q_{z}$ are in $E^{\infty}\cup E^{<}$. Thus, on all leaving arcs $e=vw$ as well as on the entering arcs $e=vw\in E^{\infty}$ , we have $y_{e}=\nu_{e}\lambda_{w}$, while $y_{e}=0$ on $e\in E^{<}$. Now, consider the flow $x^{\prime}$ on arcs crossing the cut. On all arcs leaving $Q_{z}$ we have either $0=x_{e}^{\prime}<y_{e}$ or $x^{\prime}_{e}\leq\ell^{\prime}_{w}\nu_{e}=\lambda_{w}\nu_{e}=y_{e}$. On the other hand, on the arcs entering $Q_{z}$ we have $x_{e}^{\prime}\geq 0=y_{e}$ for $e\in E^{<}$ and $x^{\prime}_{e}=\ell^{\prime}_{w}\nu_{e}=\lambda_{w}\nu_{e}=y_{e}$ for $e\in E^{\infty}$, by the WSS property, using that every such arc is always resetting and hence lies on an active $s$-$t$-path. Since $x^{\prime}$ and $y$ are both $s$-$t$-flows of value $u_{0}$ and hence have the same net flow across $Q_{z}$, all inequalities must hold with equality. In particular $x_{e}^{\prime}=0$ for every $e\in E^{<}$ crossing $Q_{z}$, and every arc $e\in E^{>}$ that crosses $Q_{z}$ is flow carrying in $x^{\prime}$. In the latter case, $v$ and $w$ both certainly lie on an active $s$-$t$-path, and so $\ell^{\prime}_{w}>\ell^{\prime}_{v}$ by the WSS property, meaning that (since we are at a point of differentiability) $e\in E^{*}_{\theta}$. Thus, we showed the properties we want for all arcs crossing some $Q_{z}$. But every arc in $E^{>}$ and $E^{<}$ crosses $Q_{z}$ for some choice of $z$, and so the claim follows.
@@
    Dual feasibility has already been observed in Section 4. Primal feasibility follows straightforwardly from (i):
@@
    • For arcs $e=vw\in E^{>}\cup E^{\infty}$ we have that $e\in E^{*}_{\theta}$ and so $x_{e}^{\prime}=\ell_{w}^{\prime}\nu_{e}=\lambda_{w}\nu_{e}=\hat{\nu}_{e}$ (using that $x^{\prime}_{e}>0$ by Section 2.2).
@@
Immediate, since $(d(\theta),p(\theta))$ is a feasible solution to the dual by Section 4, and $\Phi(\theta)$ is precisely the objective value of this dual solution. ∎
@@
We claim that $\Phi^{\prime}\geq-\sum_{e=vw\in\tilde{E}}\hat{\nu}_{e}\cdot\left(\frac{\ell^{\prime}_{w}}{\lambda_{w}}-\frac{\ell^{\prime}_{v}}{\lambda_{v}}\right)$. This follows from comparing terms. For arcs in $E^{<}\setminus E^{*}_{\theta}$ and $E^{=}\setminus E^{*}_{\theta}$, the contributions to both terms are zero. For the arc $ts$ and arcs in $^{\infty}, E>E$$^{>}\cap E^{\prime}_{\theta}$ and $E^{=}\cap E^{*}_{\theta}$, the contributions to both terms are identical.
For $e=vw\in E^{>}\setminus E^{\prime}_{\theta}$, we have
$\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{v}}\right)-\nu_{e}s^{\prime}_{e}=\hat{\nu}_{e}\ell^{\prime}_{v}\left(\tfrac{1}{\lambda_{v}}-\tfrac{1}{\lambda_{w}}\right)\geq 0.$
@@
$\nu_{e}q^{\prime}_{e}=-\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{w}}\right)\geq-\hat{\nu}_{e}\left(\tfrac{\ell^{\prime}_{w}}{\lambda_{w}}-\tfrac{\ell^{\prime}_{v}}{\lambda_{v}}\right).$
@@
Since $\dot{x}$ is a circulation and by (i), it follows that the incoming flow on arcs in $\tilde{E}$ is not smaller than the outgoing flow on arcs in $\tilde{E}$, i.e.,
[…]
@@
[…]
@@
[…]
By Section 4, $\Phi^{\prime}\geq 0$.
@@
We claim that any active arc $e=vw\in E^{\prime}_{\theta}$ satisfies $\ell^{\prime}_{w}/\lambda_{w}\geq\ell^{\prime}_{v}/\lambda_{v}$. For if not, then choosing any $z\in(\ell^{\prime}_{w}/\lambda_{w},\ell^{\prime}_{v}/\lambda_{v})$ and applying Section 4, $e\in\tilde{E}$, which we have just shown is impossible. Since $\ell^{\prime}_{s}=\lambda_{s}=1$, and every node is reachable from $s$ via an active path, $\ell^{\prime}_{v}\geq\lambda_{v}$ for all $v$.
@@
By Section 4, $\Phi$ is nondecreasing, and strictly increasing at time $\theta$ unless the WSS property is satisfied at time $\theta$. Since the WSS property implies optimality by Section 4, we deduce that $\Phi$ must be strictly increasing until the moment that $\Phi(\theta)=\mathsf{OPT}$, at which point the WSS property is satisfied for all larger times.
To bound this time (and see that it is finite), note that $\Phi^{\prime}(\theta)$ takes on only a finite set of values depending only on the network, since its value is fully determined by the configuration $(E^{\prime}_{\theta},E^{*}_{\theta})$. More concretely, for every possible valid configuration of active and resetting edges the $\ell^{\prime}$ values are determined by the thin flow conditions and thus only depend on the capacities. So we can choose $C_{1}$ depending only on the network such that $1/C_{1}\leq\Phi^{\prime}(\theta)$ for all points of differentiability $\theta$ where $\Phi^{\prime}(\theta)>0$. So $\Phi(\theta)=\mathsf{OPT}$ for all $\theta\geq C_{1}(\mathsf{OPT}-\Phi(0))$, as desired. ∎
We remark that it follows from this that if $\Phi(\theta)=\mathsf{OPT}$, then $\ell$ satisfies the WSS property. So while we stated Section 4 (ii) as an implication, the WSS property is in fact a characterization of optimality.
@@
We will need the following technical lemma.
Lemma 4.12.
Assume $\ell$ has the WSS property at time $\theta$. Then every $v\in V$ is reachable in $(V,E^{=})$ from an active node.
Proof.
Let $S$ be the set of inactive nodes that cannot be reached in $(V,E^{=})$ from an active node, and suppose for a contradiction that $S\neq\emptyset$. There are no arcs in $E^{=}$ entering $S$ by definition. By Section 4 and Section 2.2, there can be no arcs of $E^{>}$ entering or leaving $S$ (since any such arcs have $x^{\prime}_{e}(\theta)>0$, and hence have active endpoints). Similarly, by Section 2.2 no arc of $E^{\infty}$ enters or leaves $S$.
Let $y$ be a flow so that $(\lambda,y)$ is a full solution to the thin flow equations for configuration $(E,E^{\infty})$. Recall that $\lambda$ is the thin flow direction and thus $(\lambda,y)$ is a maximal solution. We will show that this leads to a contradiction. Let $\bar{\lambda}=\lambda+\varepsilon\chi(S)$, for $\varepsilon>0$ small enough that $\bar{\lambda}_{z}\geq\bar{\lambda}_{v}$ for all $zv\in\delta^{-}(S)\cup\delta^{+}(S)$. It is possible to choose such an $\varepsilon$, since $S$ has no incoming arcs in $E^{=}\cup E^{>}\cup E^{\infty}$ and no outgoing arcs in $E^{>}\cup E^{\infty}$. Since $y_{e}=0$ on all arcs entering $S$, leaving $S$, or within $S$, it is easy to check that $\bar{\lambda}$ still satisfies (2), a contradiction. ∎
@@
Consider any point of differentiability $\theta$ for which $\Phi(\theta)=\mathsf{OPT}$. Then for any $v$, $\gamma_{v}(\theta)=0$ if and only if $\ell^{\prime}_{v}(\theta)=\lambda{{}_{v}}$.
@@
Since $\Phi(\theta)=\mathsf{OPT}$, $\Phi^{\prime}(\theta)=0$, and so by Section 4 the WSS property is satisfied.
Consider any active node $v\neq s$, and let $e=uv$ be any active arc entering $v$. Then $\ell^{\prime}_{v}(\theta)=\lambda_{v}$ and $\ell^{\prime}_{u}(\theta)=\lambda_{u}$ by the WSS property. Since $e$ is active and $\theta$ is a point of differentiability, either $\ell^{\prime}_{v}(\theta)=\ell^{\prime}_{u}(\theta)$, or $e\in E^{*}_{\theta}$ (meaning $x^{\prime}_{e}(\theta)>0)$ and so $e\notin E^{<}$ by Section 4. Either way, $e\in E\setminus E^{<}$. Hence, there is an active $s$-$v$-path of arcs in $E\setminus E^{<}$, i.e., a path without slack. It follows that $\gamma_{v}(\theta)=0$ for all active nodes $v$.
Consider now any inactive node $v$. Let $P$ be a path of active arcs to $v$ starting from an active node $u$, and with all other nodes of $P$ being inactive. If there are multiple choices for $P$, choose one with the fewest number of arcs in $E^{<}$. No arc of $P$ is resetting (since (end) nodes of resetting arcs are always active). By Section 4, this further implies that no arc of $P$ is in $E^{>}\cup E^{\infty}$, so $\lambda_{u}\geq\lambda_{v}$. Since $\theta$ is a point of differentiability, it also follows that $\ell^{\prime}_{u}(\theta)=\ell^{\prime}_{v}(\theta)$.
If $\gamma_{v}(\theta)=0$, then $P$ contains only arcs of $E^{=}$ (since there must be an active path from $s$ to $v$ in $(V,E\setminus E^{<})$, and hence such a path from the set of active nodes, and $P\subseteq(E^{<}\cup E^{=}$)). So $\ell^{\prime}_{v}(\theta)=\lambda_{v}$. Conversely, if $\ell^{\prime}_{v}(\theta)=\lambda_{v}$, so that $\lambda_{u}=\ell^{\prime}_{u}(\theta)=\ell^{\prime}_{v}(\theta)=\lambda_{v}$, then $P\subseteq E^{=}$, and $\gamma_{v}(\theta)=\gamma_{u}(\theta)=0$. ∎
@@
We show that for any $v\in V\setminus S$, $\gamma^{\prime}_{v}(\theta)\leq-\Delta^{-1}<0$ (or if $\Delta=0$, then $S=V$). This suffices: after time $T_{2}=\Delta\cdot\max_{v\in V}\gamma_{v}(T_{1})$ has passed beyond time $T_{1}$, the slack $\gamma_{v}(\theta)$ has decreased to 0 for every $v\in V$, and hence all nodes are in $S$.
Consider any $v$ with $\gamma_{v}(\theta)>0$; note that $v\notin S$ by Section 4. By Section 4, $v$ is reachable from $S$ in $(V,E^{=})$. So choose $P_{1}$ to be a $u$-$v$-path in $E^{=}\setminus E[S]$ with $u\in S$ chosen to have minimum possible slack, and moreover, where this remains true for some small interval of time. That is, $\sum_{e\in P_{1}}s_{e}(\xi)=\gamma_{v}(\xi)$ for all $\xi\in[\theta,\theta+\varepsilon)$ for some small $\varepsilon>0$. This is possible given that $\theta$ is a point of differentiability; we can choose $\varepsilon$ so that all slacks are linear in this interval.
Let $P_{2}$ be a $w$-$v$ path in $E^{\prime}_{\theta}\setminus E[S]$ with $w\in S$. We have that $\lambda_{v}=\lambda_{u}=\ell^{\prime}_{u}$ and $\lambda_{v}<\lambda_{w}=\ell^{\prime}_{w}=\ell^{\prime}_{v}$. The inequality follows since $P_{2}$ must use at least one arc from $E^{<}$, or otherwise $P_{2}$ would be a no slack path in $E\setminus E^{<}$. This observation further implies that if $\Delta=0$, meaning $E^{<}=\emptyset$, then $S=V$.